OpenStudy (asadkarim7):

Two fair dice are rolled . Find the probability that the total on the two dice is 8. Given that neither shows a five.

OpenStudy (anonymous):

the probability is that it could be a 3

OpenStudy (anonymous):

neither meaning none show 5

OpenStudy (anonymous):

what do you mean by neither shows a 5 ? does the 5 is excluded from the sample space ?

OpenStudy (anonymous):

{(4,4),(6,2),(2,6)} probility is 3

OpenStudy (asadkarim7):

Given that neither die shows a five.

OpenStudy (anonymous):

cool then 3/36 = 1/12

OpenStudy (asadkarim7):

i got that answer tooo buh , unfortunately the answer is 1/4

OpenStudy (anonymous):

then the question is not properly set I guess..

OpenStudy (asadkarim7):

not possible its from the cambridge book , it cant be

OpenStudy (anonymous):

how?

OpenStudy (across):

Shouldn't it be: (6,2),(2,6),(4,4),(4,4)?

OpenStudy (asadkarim7):

then it would be 4/ 36 so 1/9 . answer is still wrong

OpenStudy (across):

I'm also tempted to believe there must be a typo somewhere... or there's something we don't know. Rolling the first dice and having it land on either 2, 4 or 6 gives a probability of 1/2. The second dice then has a 1/6 probability of landing in such a way that it sums to 8 with the previous dice. This gives 1/12. -.-

OpenStudy (anonymous):

@across: (4,4) and (4,4) both are same events..

OpenStudy (asadkarim7):

ooh Sweet , i got itt . :) yupiie there are 3 outcomes (4|4) or ( 6|2)or (2|6) so Each has probabilty 1/12 1/12 + 1/12+1/12 therefore 3/12 1/4

OpenStudy (across):

Landing a (4,4) has a probability of 1/12!?

OpenStudy (anonymous):

o Each has probabilty 1/12 -- how ? :O

OpenStudy (anonymous):

why don't u say(4,4) (4,4) as (6,2) (2,6)

OpenStudy (asadkarim7):

3/36 = 1/12 1/12 x 3 coz 3 outcomes 1/4

OpenStudy (anonymous):

duh ...... your probability skills are just wonderful.. congratulation.

OpenStudy (across):

1/12 is already the probability of landing a dice sum of 8 without taking 5 into consideration. Then you're multiplying this probability by 3? Are you casting the dice thrice?

OpenStudy (asadkarim7):

Umm i dont knnow